Find the sum of the x and y intercept of any tangent line to the curve √x + √y = √k
1
Expert's answer
2020-04-20T13:37:35-0400
"\sqrt{x}+\sqrt{y}=\sqrt{k}"

"{d\over dx}(\sqrt{x})+{d\over dx}(\sqrt{y})={d\over dx}(\sqrt{k})"

"{1 \over 2\sqrt{x}}+{1 \over 2\sqrt{y}}y'=0"

"y'=-{\sqrt{y} \over \sqrt{x}}"

Tangent line


"y-y_0=-{\sqrt{y_0} \over \sqrt{x_0}}(x-x_0)"

"y=-{\sqrt{y_0} \over \sqrt{x_0}}x+\sqrt{x_0}\sqrt{y_0}+y_0""y=-{\sqrt{y_0} \over \sqrt{x_0}}x+\sqrt{y_0}(\sqrt{x_0}+\sqrt{y_0})"

"y=-{\sqrt{y_0} \over \sqrt{x_0}}x+\sqrt{k}\sqrt{y_0}"

x-intercept: "y_1=0"


"0=-{\sqrt{y_0} \over \sqrt{x_0}}x_1+\sqrt{k}\sqrt{y_0}"

"x_1=\sqrt{k}\sqrt{x_0}"

yintercept: "x_2=0"


"y_2=-{\sqrt{y_0} \over \sqrt{x_0}}(0)+\sqrt{k}\sqrt{y_0}"

"y_2=\sqrt{k}\sqrt{y_0}"

 The sum of the x and y intercept of any tangent line to the curve is


"x_1+y_2=\sqrt{k}\sqrt{x_0}+\sqrt{k}\sqrt{y_0}=\sqrt{k}(\sqrt{x_0}+\sqrt{y_0})="

"=\sqrt{k}\sqrt{k}=k"

The sum of the x and y intercept of any tangent line to the curve is equal to "k."

At point "(0, \sqrt{k})" the tangent line is "x=0."

At point "(\sqrt{k},0)" the tangent line is "y=0."